The difference between two numbers is 12, and four times the
Langer number exceeds three times the smaller number by 90,what is the larger number?

Answer by Maths68(1474)

Let
Larger number = x
Smaller number = y
Given
======
The difference between two numbers is 12
x-y=12
x=12+y..................(1)
Four times the larger number exceeds three times the smaller number by 90
4x=3y+90...........(2)
Put the value of x from (1) to (2)
4x=3y+90
4(12+y)=3y+90...........(2)
48+4y=3y+90
4y-3y=90-48
y=42
Put the value of y in (1)
x=12+y
x=12+42
x=54

Larger Number is 54